OpenStudy (canimcan):

I am thinking about this integral (x^2)*(sqrt (x^2+16))🔍 Can I rationalize it? *** One more question about Ostogradski formulas -does it exist his formula for integrals such is (ax^3+bx^2+cx+d)(sqrt...)+λ/(sqrt..)...?

OpenStudy (freckles):

Did you try a trig sub?

OpenStudy (karim728):

first one works great using trig sub

OpenStudy (canimcan):

Yes, I take x=4tgt but I can't finish .😕

OpenStudy (holsteremission):

\[\begin{align*}\int x^2\sqrt{x^2+16}\,\mathrm{d}x&=\int (4\tan t)^2\sqrt{(4\tan t)^2+4^2}\,\mathrm{d}t\\[1ex] &=16\sqrt{16}\int\tan^2t\sqrt{\tan^2t+1}\,\mathrm{d}t\\[1ex] &=64\int \tan^2t\sec t\,\mathrm{d}t&(*) \end{align*}\](\((*)\) for valid \(t\)) Hint: If you can integrate \(\sec^3t\), the rest should be easy.

OpenStudy (freckles):

\[\int\limits x^2 \sqrt{x^2+16} dx=\int\limits (4 \tan(t))^2 \sqrt{(4 \tan(t))^2+4^2} 4 \sec^2(t) dt \\ \text{ since } dx \neq dt \text{ but } dx=4 \sec^2(t) dt \\ \]

OpenStudy (holsteremission):

Whoops, my mistake. The last line should then be \[256\int\tan^2t\sec^3t\,\mathrm{d}t\]Thanks @freckles